OpenStudy (otonogold):

Simplify 3^1/2 * 3^1/2. Can you show me steps on how to simplify this?

OpenStudy (potatoes.ramu):

that's basically \[\sqrt{3}*\sqrt{3}\]

OpenStudy (potatoes.ramu):

Or 3^(1/2+1/2)

OpenStudy (potatoes.ramu):

Which is 3

OpenStudy (otonogold):

Im confused.

OpenStudy (otonogold):

@jdoe0001 Can you explain this for me?

OpenStudy (potatoes.ramu):

There's a formula for this

OpenStudy (potatoes.ramu):

It's a^m*a^n=a^(m+n)

OpenStudy (jdoe0001):

\(\Large \bf {a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\\ \quad \\ \quad \\ 3^{\frac{{\color{blue} 1}}{{\color{red} 2}}}\cdot 3^{\frac{{\color{blue} 1}}{{\color{red} 2}}}\implies \sqrt[{\color{red} 2}]{3^{\color{blue} 1}}\cdot \sqrt[{\color{red} 2}]{3^{\color{blue} 1}}}\)

OpenStudy (jdoe0001):

\(\bf \implies \sqrt{3}\cdot \sqrt{3}\)

OpenStudy (otonogold):

then being 3 ? as the answer...

OpenStudy (jdoe0001):

yeap

OpenStudy (otonogold):

Thank you (:

OpenStudy (jdoe0001):

yw